OpenStudy (drewbooks):

What is the simplest form of the expression below? 3a / b × b^2 / 12c

OpenStudy (drewbooks):

Attached to this post is the list of answers. I don't know if I'm solving this correctly, but the answer I'm coming up with is the first. Is that right?

OpenStudy (theopenstudyowl):

d

OpenStudy (drewbooks):

Oh, really? How do you come up with that?

OpenStudy (theopenstudyowl):

medal me and I will explain..

OpenStudy (drewbooks):

Well, I was going to give you a medal out of kindness, but okay.

OpenStudy (theopenstudyowl):

\[3a / b × b^2 / 12c=3a/1 \times b/12c = 3ab/12c = ab/4c\]

OpenStudy (theopenstudyowl):

If your able to master some of these easier ones, the harder ones won't be as hard!

OpenStudy (drewbooks):

Ah, okay. I see what I did wrong. Thanks for the help.

OpenStudy (theopenstudyowl):

np